# MIST: Mutual Information Estimation via Supervised Training

### v1

@@ A.3 Empirical Evaluation Scope of MI Estimators
In Table 3, we provide a summary of how prominent estimators were empirically evaluated. We focus our summary on evaluations of the estimator directly, rather than the performance of an estimator in a downstream tasks. Similarly, we only include the evaluations of estimators in the papers introducing the method, as we are most interested in how estimator quality is determined.
@@
KDE Time-Series 4 4 1 - 400, 2K, 4K
KSG Synthetic 4 3 3, 8 0.83,1,1.2 125, 10K, 50K
MIENF Synthetic 6 5 32, 1024 10 10K
LMI Synthetic,MNIST,PLM 3 2 5K,784,1024 1 2K,5K
@@
MINDE Synthetic 20, 3 17, 2 25, 9 1, 5 100K
VCE Synthetic,Images,Text 7,2,2 6,1,2 64,16,16 16,7,2.1 10K,10K,4K
MINE Synthetic 2 0 2, 20 2, 40 -
@@
Table 3: Summary of the empirical evaluation scope for prominent mutual information estimators. Comma-separated entries indicate values for different experimental settings (e.g., KDE has three experiments with minimum sample size considered 400, 2K, and 4K, respectively). Dashes indicate data that is not available or not specified.
@@
Early methods for mutual information estimation, such as binning, were evaluated primarily using simulated time-series data. These commonly included sine and cosine functions, as well as chaotic dynamical systems such as the Lorenz and Rossler attractors (where the task is phase reconstruction, and no ground truth is available). In the initial histogram binning work (histogram_binning), the Rossler attractor was used. For this work, we include it in the table as having one 1D distribution, which is non-Gaussian, with no ground truth MI reported (and therefore no maximum). The experiment calculated the mutual information over 65,536 points.
KDE (kde) was similarly evaluated with time-series data. KDE was evaluated on 1D synthetic sine waves, an autoregressive function of a sine and a Gaussian distribution, the Lorenz attractor, and the Rossler attractor. 400 samples were used for the sine function, and 500 samples were used for the autoregressive function. 4096 and 2048 samples were used for the Lorenz and Rossler attractors, respectively. Estimators are evaluated on the quality of the reconstruction, with no ground truth MI available. We include KDE in the table as having 4 distributions (sine, AR, Lorenz, Rossler), 4 of which are non-Gaussian, and all 1D. No ground truth MI values are present, and the samples included 400, 2K, 4K.
KSG (ksg) performed evaluation on 2D correlated Gaussians with unit variance and zero mean, testing sample sizes of 125, 250, 500, 1K, 2K, 4K, 10K, and 20K. The MI was 0.830366. Authors performed error analysis with 10K samples, and performed experiments on the gamma-exponential distribution, the ordered Weinman exponential, and the circle distribution. All of these had MI values below 1, and considered samples with 1K and 10K data points. For their higher dimension experiments, the estimator was evaluated on 3-D Gaussian distributions with a sample size of 10K in one experiment, and on m-correlated uniform Gaussians up to dimension 8 with a sample size of 50K. We include KSG in the table as being evaluated on synthetic 4 distributions (Gaussian, Weinman-exponential, gamma-exponential, and circle), 3 of which are non-Gaussian. The maximum dimension considered was 8, while the maximum MI values were 0.83, 1, and 1.2 for the 1-D Gaussians, 3-D Gaussians, and 8-D Gaussian experiments, respectively. The minimum sample size in these experiments was 125 for the 1-D Gaussian, 10K for other distributions, and 50K for larger dimensions.
MIENF (mi_through_normalizing_flows) first evaluated their estimator on 2D correlated Gaussian distributions mapped into higher-dimensional images (16x16 and 32x32), evaluated on 10K samples, and with MI values between 0 and 10. In their second experiment, authors evaluated their method on five non-compressible distributions (uniform, smoothed uniform, student (dof = 4), arcsignh (student) student with dof = 1,2). For this experiment, authors varied the dimensions of these distribution from 2 to 32, using 10K samples for each prediction. The MI was set to 10 for all distributions and dimensions.
@@
Latent-MI (LMI; lmi_estimator) was evaluated on multivariate Gaussian distributions under varying original data dimensionality and latent dimensionality learned before applying KSG. Their multivariate Gaussian experiment considered up to 10 latent dimensions and 5K original dimensions, with 2K samples. In a second experiment, the authors studied the sample requirements of their estimators. For this experiment, authors used multivariate Gaussians from dimension 1 to 50, each with one correlated dimension, and an MI of 1. Authors varied the sample size up to 4K. In a third experiment, MNIST and protein embeddings were resampled to align with a Bernoulli vector with set correlation. For MNIST, the dimension was 784, and the sample size was 5K. For the protein embeddings, the dimension was 1024 and the sample size was 4402. In both cases, datasets were generated with MI values between 0 and 1.
@@
For GM and SMILE (MINE_exp_variance_consistency_tests), evaluation was performed on two tasks. In the first task, authors evaluated their methods with 20-D Gaussians with correlation $p$, and in the second task, the authors applied a cubic operation on these Gaussians. This set-up was introduced in (mine). MI values were increased from 0 to 10 at step intervals during the critic training process. Training occurred over 20K steps, with MI increasing by 2 every 4K steps. The batch size was 64, and each batch was a new random sample. For the sake of our table, we consider the number of samples used for estimation to be the number of samples used to fit the critic for a given MI value. Since the estimator saw 4K steps of batch size 64 for each MI (2, 4, 6, 8, 10), we list the minimum samples seen as 4,000 x 64 = 256K. This work also includes consistency tests constructed by adapting MNIST images, but this setting does not carry ground-truth MI values (only the ideal ratio of estimates).
MINDE (minde) used the BMI benchmark (bmi_benchmark) at larger sample sizes (100K rather than 10K). For our table, we consider the number of unique distribution types considered; for BMI, there are 20 unique distribution tasks after excluding tasks that are higher-dimensional versions of another. The highest MI value used was 5, and the maximum dimension considered was 25. MINDE also performs image consistency experiments as in SMILE, but this setting does not have a ground truth value.
@@
VCE chen2025neural_vector_copulas evaluate estimators on five synthetic distributions with varying dependence. In this setting, MI ranges from 0 to 16, 10K samples are used, and dimensions range from 2 to 64. Then, the authors study increasing dimensionality on five distributions, with dimensions from 2 to 64 and MI values up to 14. 10K samples were used, and six distributions across the two synthetic experiments were not gaussian. Next, authors consider correlated rectangles and gaussian plates. These images have dimension 16x16 but are processed by an autoencoder prior to prediction to produce 16 dimensional embeddings. The max MI for this setting is 7 and 10K samples were used. Finally, they consider language embeddings. Data is similarly processed by an autoencoder to be 16 dimensions. The maximum MI is 2.1, and 400 samples were used.
MINE (mine) was first evaluated on 2-D and 20-D Gaussians with MI from 0 to 2 and 0 to 40, respectively. The number of samples used was not specified. In a second experiment, the authors applied a deterministic nonlinear transformation and demonstrated that their estimates are invariant to it.
@@
NWJ (nwj) introduced a variational bound for estimating the KL-divergence between two arbitrary distributions, and their experiments focused on evaluating the quality of this bound and estimation on various distributions. The authors did not evaluate this approach on mutual information prediction directly, for which the distributions compared would be a joint distribution and the product of its marginals.
InfoNCE (oord2018_cpc) was developed as a method for representation learning, with evaluation including representation learning tasks but without any evaluation of MI prediction itself.
@@ Appendix B Details of Meta-Dataset Creation
We construct a meta-dataset of synthetic distributions with known MI using the BMI library (bmi_benchmark), which generates complex yet tractable distributions via invertible transformations applied to base families with analytical MI. Each meta-sample consists of joint samples from a distribution paired with its ground-truth MI value, enabling supervised learning directly over distributions.
@@ C.1 Experiments for Normalized and Multi-Task Predictions
Figure 10: Mean MI estimates across prediction strategies on $\mathcal{M}_{\text{test}}$ ($\mathrm{dim}=9$), comparing normalization methods and coefficient configurations. “3-coef” uses the estimators $(\hat{H}_{X,\mathrm{inv}},\hat{H}_{Y,\mathrm{inv}},\hat{H}_{XY,\mathrm{inv}})$; the last two entries show predictions with $\hat{I}_{\mathrm{inv}}$, with and without the physics-informed loss $\mathcal{L}_{\mathrm{phys}}$. Values are (mean ± stddev) over three runs.
Experiments on $\mathcal{M}_{\text{test}}$ with $\mathrm{dim}=9$ are summarized in Table 10. The best results were obtained using two methods: direct MI prediction, and the four-coefficient formulation with the physics-informed loss. The comparatively lower performance of the three-coefficient model can be attributed to error propagation, since each entropy term is predicted independently and their individual errors accumulate during MI reconstruction.
In this work, we proceed with the direct MI prediction strategy for subsequent experiments. Nevertheless, the four-coefficient formulation remains a promising direction for future research, as it provides richer structural information about the underlying distributions and could facilitate the derivation of both MI and entropy values within a unified framework.
@@ C.2 Dimensionality Ablation
Figures 11 through 13 contain additional bar plots for the experiment described in the second half of Section 4.6.
@@
[…]
We give in Figures 14 through 16 a high-level summary of the general performance of all various fixed- and variable-dimensionality models for ease of comparison.
@@ C.5 Scaling Behavior
We report in Table 7 the performance of the main model (5.5M) used in our experiments compared to scaled-down models.
@@
Table 7: Average performance metrics for models of different sizes.
We give in Figure 17 a fine-grained heatmap of the MSE performance obtained by MIST and MISTQR for various sample size groups and dimensions.
@@ D.1 Self-Consistency of Quantile Regression
D.2 Self-Consistency of the Learnt Quantity
[…]
  • The independence test checks that if $X\mathrel{\perp\!\!\!\perp}Y$, then $I(X;Y)=0$.
@@ D.2 Self-Consistency of the Learnt Quantity
  • The additivity over independent samples test checks that if $(X_{1},X_{2})\mathrel{\perp\!\!\!\perp}(Y_{1},Y_{2})$, then $I(X_{1},Y_{1};X_{1},Y_{2})=I(X_{1};X_{2})+I(Y_{1},Y_{2})$
@@
Figure 19: Data Processing Test: AA is an MNIST image, and $B_{r}$ contains the top $r$ rows of image AA. This test compares $I([A,A],[B_{r},B_{{}^{\prime}r-3^{\prime}}])$ to $I([A,B_{r}])$, where AA is an original MNIST image, $B_{r}$ is the first $r$ rows of $A$, and so $B_{{}^{\prime}r-3^{\prime}}$ is a further masked version of image AA. The test produces a ratio which should be near to 1 for all values of $r$, which amounts to checking that adding a further masked version of your image $B_{r}$ does not change the predicted MI. Images are encoded with PCA fit on 10K random samples, and error bars of $1\sigma$ reported, with 100 bootstrapped samples.

### v3

@@ A.3 Empirical Evaluation Scope of MI Estimators
In Table 3, we provide a summary of how prominent estimators were empirically evaluated. We only include the evaluations of estimators in the papers introducing the method, as we are most interested in how estimator quality is determined. This summary focuses on the experiments evaluating the estimator directly, and excludes evaluations in downstream settings, where ground truth MI values are not available.
Table 3: Summary of the empirical evaluation scope of prominent mutual information estimators. Comma-separated entries indicate values for different experimental settings (e.g., KSG has four experiments). Dashes indicate data that is not available or not specified.
@@
KSG Synthetic 1,3,1,1 0,3,0,0 2,1,3,8 0.83,1,1,1.2 125,1K,10K,50K
KDE Time-Series 4 4 1 - 400
MIENF Synthetic 1,5 0,5 1024,32 10,10 10K,10K
LMI Synthetic,Images,PLM 1,1,1,1 0,0,1,1 5K,50,784,1024 2,1,1,1 2K,100,5K,4.4K
@@
MINDE Synthetic 19,3 16,2 25,9 1,5 100K,-
VCE Synthetic,Images,Text 5,5,2,2 4,3,1,2 64,64,16,16 16,14,7,2.1 10K,10K,10K,4K
MINE Synthetic 2,3 0,3 20,2 40,- -,-
@@
1 Using slicing, which takes an expectation over random 1D projections
@@
Early methods for mutual information estimation, such as binning, were evaluated primarily using simulated time-series data. These commonly included sine and cosine functions, as well as chaotic dynamical systems such as the Lorenz and Rossler attractors (where the task is phase reconstruction, and no ground truth is available). In the initial histogram binning work (Fraser and Swinney, 1986), the Rossler attractor was used. For this work, we include it in the table as having one 1D distribution, which is non-Gaussian, with no ground truth MI reported (and therefore no maximum). The experiment calculated the mutual information over 65,536 points. KDE (Moon et al., 1995) was similarly evaluated with time-series data. KDE was evaluated on 1D synthetic sine waves, an autoregressive function of a sine and a Gaussian distribution, the Lorenz attractor, and the Rossler attractor. 400 samples were used for the sine function, and 500 samples were used for the autoregressive function. 4096 and 2048 samples were used for the Lorenz and Rossler attractors, respectively. Estimators are evaluated on the quality of the reconstruction, with no ground truth MI available. KSG (Kraskov et al., 2004) performed evaluation on 2D correlated Gaussians with unit variance and zero mean, testing sample sizes of 125, 250, 500, 1K, 2K, 4K, 10K, and 20K. The MI was 0.830366. Authors performed error analysis with 10K samples, and performed experiments on the gamma-exponential distribution, the ordered Weinman exponential, and the circle distribution. All of these had MI values below 1, and considered samples with 1K and 10K data points. For their higher dimension experiments, the estimator was evaluated on 3-D Gaussian distributions with a sample size of 10K in one experiment, and on m-correlated uniform Gaussians up to dimension 8 with a sample size of 50K. We include KSG in the table as being evaluated on synthetic 4 distributions (Gaussian, Weinman-exponential, gamma-exponential, and circle), 3 of which are non-Gaussian. The maximum dimension considered was 8, while the maximum MI values were 0.83, 1, and 1.2 for the 1-D Gaussians, 3-D Gaussians, and 8-D Gaussian experiments, respectively. The minimum sample size in these experiments was 125 for the 1-D Gaussian, 10K for other distributions, and 50K for larger dimensions. MIENF (Butakov et al., 2024) first evaluated their estimator on 2D correlated Gaussian distributions mapped into higher-dimensional images (16x16 and 32x32), evaluated on 10K samples, and with MI values between 0 and 10. In their second experiment, authors evaluated their method on five non-compressible distributions (uniform, smoothed uniform, student (dof = 4), arcsignh (student) student with dof = 1,2). For this experiment, authors varied the dimensions of these distribution from 2 to 32, using 10K samples for each prediction. The MI was set to 10 for all distributions and dimensions.
@@
Latent-MI (LMI; Gowri et al., 2024) was first evaluated on multivariate Gaussian distributions under varying original data dimensionality and latent dimensionality (the dimensionality of the KSG input). Their multivariate Gaussian experiment considered up to 10 latent dimensions and 5K original dimensions, with 2K samples. The MI ranged from 0 to 2 bits. In a second experiment, the authors studied the sample requirements of their estimators. For this experiment, authors used multivariate Gaussians from dimension 1 to 50, each with one correlated dimension, and an MI of 1. Authors varied the sample size from 100 to 10K. In a third experiment, MNIST and protein embeddings were resampled to align with a Bernoulli vector with set correlation. We consider the MIST and protein embedding experiments separately. For MNIST, the dimension was 784, and the sample size was 5K, and the dataset was generated with MI values between 0 and 1. For the protein embeddings, the dimension was 1024, the sample size was 4402, and the dataset was generated with MI values between 0 and 1.
@@
For GM and SMILE (Song and Ermon, 2019), evaluation was initially performed two distributions. The first distribution was a 20-D Gaussian with correlation $p$ (borrowing from a setup in (Belghazi et al., 2018)), and the second was this distribution with a cubic operation applied. Authors follow a setup from (Poole et al., 2019), in which MI values were increased from 0 to 10 at step intervals during the critic training process. Training occurred over 20K steps, with MI increasing by 2 every 4K steps. The batch size was 64, and each batch was a new random sample. For the sake of our table, we consider the number of samples used for estimation to be the number of samples used to fit the critic for a given MI value. Since the estimator saw 4K steps of batch size 64 for each MI (2, 4, 6, 8, 10), we list the minimum samples seen as 4,000 x 64 = 256K. In a second set of experiments, both methods are evaluated for consistency tests with pairs of MNIST and CIFAR images. Because these do not have ground truth MI values, we exclude them from the table. In this work, a flow-based GM model is used for the first experiment, and a VAE-based GM model is used for the consistency experiment.
MINDE (Franzese et al., 2023) used the BMI benchmark (Czyż et al., 2023) to evaluate their model. BMI contains 40 tasks. We count any unique combination of distribution and transform (including degrees of freedom and noise levels) to be a unique distribution type, excluding tasks that only differ by the dimensions or by correlation strength. Among the 40 tasks, there are 19 of these unique distributions, and 16 of these are non-normal. In their first experiment, they evaluate on these distributions with 100K samples, with MI ranging from 0.2 to 1. In a second experiment, they evaluate on 3x3 multivariate normal distributions, along with a half-cube and spiral transformation of them. Their experiment increases MI from 0 to 5. We count this experiment as having 3 unique distribution types, 2 of which are non-normal. The sample size is not directly reported. In a third experiment, they repeat the consistency tests from SMILE (Song and Ermon, 2019) on MNIST only, which we exclude because there is no ground truth MI.
@@
VCE Chen et al. (2025) first evaluate estimators on five synthetic distributions, under varying dependence. In this setting, MI ranges from 0 to 16, 10K samples are used, and dimensions range from 2 to 64. In a second experiment, the authors study increasing dimensionality on five distributions (three non-normal), with dimensions from 2 to 64 and MI values up to 14. 10K samples were used. In a third experiment, authors consider correlated rectangles and gaussian plates. These images have dimension 16x16 but are processed by an autoencoder prior to prediction to produce 16 dimensional embeddings. The embedding portion is considered a pre-processing step, rather than part of the estimation method (as in LMI), and we therefore report the embedding dimension. This is consistent with other similar experiments, in which lower-dimensional data is scaled up to higher dimensions prior to evaluation, for which we correspondingly select the scaled dimensionality. The dimension reported is the dimension of data provided to the estimator. We consider this experiment to have 2 distributions, 1 of which is non normal, and dimension of 16. The maximum MI value is 7, and 10K samples used. In a fourth experiment, they consider language embeddings from Llama-3 13B and Bert. We consider this to be 2 non-normal distributions. Data is similarly processed by an autoencoder to be 16 dimensions. The underlying MI is estimated with resampling, and the maximum reported value is 2.1. 4K samples were used.
@@
MINE (Belghazi et al., 2018) was first evaluated on 2-D and 20-D Gaussians with MI from 0 to 2 and 0 to 40, respectively. The number of samples used was not specified. In a second experiment, the authors applied 3 deterministic nonlinear transformations to a normal distribution and demonstrated that their estimates are invariant to it. MINE then performs several downstream application experiments. NWJ (Nguyen et al., 2010) introduced a variational bound for estimating the KL-divergence between two arbitrary distributions, and their experiments focused on evaluating the quality of this bound and estimation on various distributions. The authors did not evaluate this approach on mutual information prediction directly, for which the distributions compared would be a joint distribution and the product of its marginals. InfoNCE (Oord et al., 2018) was developed as a method for representation learning, with evaluation including representation learning tasks but without any evaluation of MI prediction itself.
FMMI (Butakov et al., 2025) empirically evaluate their method on four distribution types (correlated normal, half-cube correlated normal, correlated uniform, smoothed uniform), with MI ranging from 0 to 80 nats. We convert nats (natural logarithm, base $e$) to bits (base 2), by dividing the nats value by $\log_{e}(2)$, and round to the nearest whole number (115). In this experiment, the training size was 100K samples, and the test set was 1K samples. The dimensions ranged from 0 to 160 dimensions. FMMI is also applied to predict O-information in fMRI data.
InfoNet (Hu et al., 2024) perform their first experiment evaluating test-time efficiency on GMM distributions of various lengths, using sequence of 200, 500, 1000, 2000, and 5000 samples. Since no additional details about the estimation are included (maximum MI, dimension, number of samples), we exclude it from the table. In their second experiment (the first we report in the table), they benchmark MI estimation on 1D gaussian distributions with varying correlation. The MI ranges from 0 to 0.9, and 2000 samples are used. In the third experiment (the second we report in the table), the authors evaluate their method on 1D GMMs with multiple components (up to 20, which we consider as 20 distinct distribution types, 19 of which are non-gaussian), using 10 MI levels from 0 to 0.9. Authors use 2K samples. In their fourth experiment (the third we report in the table), the authors evaluate order accuracy on 1D GMMs with components from 1 to 10 (we count this as 10 distinct distributions, 9 of which are non-normal). The number of samples and maximum MI value are not reported. In their fifth experiment (the fourth we report in the table), they evaluate an application in high-dimensional independence testing, using slicing to scale their 1D-trained model. Their experiment tests three structures of correlation in gaussian distributions, and we report this as 3 distribution types, 0 of which are non-gaussian. They evaluate on 16 and 128 dimensions, and the number of samples varied from 50 to 950. In their sixth experiment, they evaluate their approach using MI to perform object segmentation in video data. Because this experiment is a downstream application without ground-truth MI values, we exclude it from the table.
In this work, MIST, we use the BMI library to generate a range of 23 distribution types (16 in $\mathcal{M}_{\text{train}}$, 7 in $\mathcal{M}_{\text{test,OoMD}}$). Of these distributions, 19 are non-normal. The maximum dimension we use is 32, and we test MI values up to 40. The minimum sample size in our experiments is 10. Our downstream experiments in consistency tests and feature selection are excluded from this table, whose focus is evaluating MI estimation directly.
@@ Appendix B Details of Meta-Dataset Creation
We construct a meta-dataset of synthetic distributions with known MI using the BMI library (Czyż et al., 2023), which generates complex yet tractable distributions via invertible transformations applied to base families with analytical MI. Each meta-sample consists of joint samples from a distribution paired with its ground-truth MI value, enabling supervised learning directly over distributions.
@@ C.1 Experiments for Normalized and Multi-Task Predictions
Table 5: Mean MI estimates across prediction strategies on $\mathcal{M}_{\text{test}}$ ($\mathrm{dim}=9$), comparing normalization methods and coefficient configurations. “3-coef” uses the estimators $(\hat{H}_{X,\mathrm{inv}},\hat{H}_{Y,\mathrm{inv}},\hat{H}_{XY,\mathrm{inv}})$; the last two entries show predictions with $\hat{I}_{\mathrm{inv}}$, with and without the physics-informed loss $\mathcal{L}_{\mathrm{phys}}$. Values are (mean ± stddev) over three runs.
@@
Experiments on $\mathcal{M}_{\text{test}}$ with $\mathrm{dim}=9$ are summarized in Table 9. The best results were obtained using two methods: direct MI prediction, and the four-coefficient formulation with the physics-informed loss. The comparatively lower performance of the three-coefficient model can be attributed to error propagation, since each entropy term is predicted independently and their individual errors accumulate during MI reconstruction. In this work, we proceed with the direct MI prediction strategy for subsequent experiments. Nevertheless, the four-coefficient formulation remains a promising direction for future research, as it provides richer structural information about the underlying distributions and could facilitate the derivation of both MI and entropy values within a unified framework.
@@ C.2 Dimensionality Ablation
Figures 10 and  11 contain additional bar plots for the experiment described in the second half of Section 4.7. We give in Figures 12 through 14 a high-level summary of the general performance of all various fixed- and variable-dimensionality models for ease of comparison.
@@ C.5 Scaling Behavior
We report in Table 8 the performance of the main model (5.5M) used in our experiments compared to scaled-down models.We give in Figure 15 a fine-grained heatmap of the MSE performance obtained by MIST and MISTQR for various sample size groups and dimensions.
Table 8: Average performance metrics for models of different sizes.
@@ D.2 Self-Consistency of the Learned Quantity
  • The independence test checks that if $X\perp\!\!\!\!\perp Y$, then $I(X;Y)=0$.
@@
  • The additivity over independent samples test checks that if $(X_{1},X_{2})\perp\!\!\!\!\perp(Y_{1},Y_{2})$, then $I(X_{1},Y_{1};X_{1},Y_{2})=I(X_{1};X_{2})+I(Y_{1},Y_{2})$
@@
Figure 17: Data Processing Test: AA is an MNIST image, and $B_{r}$ contains the top $r$ rows of image AA. This test compares $I([A,A],[B_{r},B_{{}^{\prime}r-3^{\prime}}])$ to $I([A,B_{r}])$, where $A$ is an original MNIST image, $B_{r}$ is the first $r$ rows of $A$, and so $B_{{}^{\prime}r-3^{\prime}}$ is a further masked version of image AA. The test produces a ratio which should be near to 1 for all values of $r$, which amounts to checking that adding a further masked version of your image $B_{r}$ does not change the predicted MI. Images are encoded with PCA fit on 10K random samples, and error bars of $1\sigma$ reported, with 100 bootstrapped samples.
@@
Appendix E Feature Selection
In this section, we explore using our estimators for a simple feature selection problem in which feature weights are learned by maximizing the mutual information with an outcome Y. We define $X\in\mathbb{R}^{n\times d}$ as a random normal matrix with $d$ features and $n$ samples. Next, we define a set of weights for these features, each bounded between 0 and 1, $w\in[0,1]^{d\times 1}$. The label Y is given by $Y_{\text{true}}=Xw^{*}+\epsilon$, where $Y\in\mathbb{R}^{n\times 1}$ is the matrix multiplication of $X$ and pre-set masking weights $w^{*}$ plus added normal noise $\epsilon$. For a candidate set of weights, the predicted label is $Y^{\prime}=Xw^{\prime}\in\mathbb{R}^{n\times 1}$, produced by the candidate weights $w^{\prime}$. The best weights are chosen as $w_{\text{best}}=\operatorname*{arg\,max}_{w^{\prime}}\text{MIST}(Y_{\text{true}},Y^{\prime})$, which maximizes the mutual information with the true label.
The goal of this task is to recover the true weights ($w^{*}$), which correctly assign a value of 1 for the two informative features and 0 otherwise. We set the mean and variance of $X$ to 0 and 1.0 respectively, and the normal noise applied to $Y_{true}$ has 0 mean and variance of either 0.3, 0.5, or 1.0. We generate 32 features and randomly select 2 features to be informative. We use MIST and MINE estimators, and perform gradient ascent with the Adam optimizer and learning rate of 0.01 for 500 steps. For MINE, we consider two settings. In the first, we continually train one MINE estimator, allowing 15 steps of updates with each prediction. In the second setting, we train a new MINE estimator for 200 steps for each prediction. In both cases, we follow the parameters in MINE’s repository, using Adam for the optimizer and a learning rate of 0.001. We used the biased version loss and set the critic hidden dimension to 64. Weights are initialized with a small constant and bounded between 0 and 1.
In Figure 19, we plot the ROC curves of the learned features, which includes 3 noise values (0.3, 0.7, and 1.0) and 20 random seeds each. We find that even with a very limited number of samples, MIST estimates provide an accurate and stable learning signal. Training MINE to convergence at each prediction achieves better performance than continually MINE training, but remains significantly less effective than using MIST estimates. In Figure 20, we plot the top-k accuracy according to the number of samples provided. Accuracy is reported as the proportion of the two highest weighted features that are truly informative. While the number of informative features is rarely known ahead of time, we provide this view of the results to better compare the method performance at each sample size. We find that MIST-guided selection is more accurate and lower variance on average, achieving perfect accuracy at 100 samples. In Figure 21, we show the feature weights through training of two randomly chosen seeds with noise variance set to 1.0.
Figure 19: ROC curves of feature selection learned with MINE and MIST estimates. Results are computed with 3 noise levels and 20 random seeds, with 1 standard deviation shaded. We find that MIST-guided feature selecting is more accurate and lower variance across settings.
Figure 20: Comparison of the accuracy of features selected via gradient ascent on MINE predictions (either trained continually and trained to convergence) and MIST predictions. Accuracy is calculated as the proportion of the highest two weights which are the informative features. In other words, how often are top 2 weights assigned to the two correlated features? We find that MIST-guided feature selection is more accurate and lower variance.
MINE (Continual)
MINE (Convergence)
MIST
Figure 21: Comparison of feature selection guided by MINE with continual training (left), MINE with convergence training (middle), and MIST (right). Plotted are the feature weights throughout training on two randomly selected seeds of data (rows). We find MIST-guided training to be very effective at differentiating correlated features.